OpenStudy (anonymous):

the function f is defined for all real numbers x by f(x) = ax^2 + bx + c where a, b and c are constants and a is negative, in the xy-plane, the x coordinate of the vertex of the parabola is y=f(x) is -1. if t is a number for which f(t) > f(0), which of the following (any or all) must be true? 1: -2 < t < 0 2: f(t) < f(-2) 3: f(t) > f(1) please help!

myininaya (myininaya):

we know this parabola is oven downward because a is negative and the vertex is (-1,k)

myininaya (myininaya):

f(t)>f(0) ....

myininaya (myininaya):

f(t)=a(t+1)^2+k and f(0)=a(0+1)^2+k=a+k and f(t)>f(0) so we have a(t+1)^2+k=a+k

myininaya (myininaya):

solve for t

myininaya (myininaya):

so -2<t<0

OpenStudy (anonymous):

so are you solving as a system here?

myininaya (myininaya):

i solve that equation by using remedial algebra

myininaya (myininaya):

subtract k on both sides

myininaya (myininaya):

divide by a on both sides

OpenStudy (anonymous):

so (t+1)^2 = 1

myininaya (myininaya):

yes

myininaya (myininaya):

square root both sides

myininaya (myininaya):

you get t+1=+ or - 1

myininaya (myininaya):

t=-1-1,1-1=-2,0

OpenStudy (anonymous):

ok, i think i understand that. also, it says option 3 is true, can you help with that?

myininaya (myininaya):

lets see we gots to think I can't believe I could actually about the first one lol

myininaya (myininaya):

i have a question where i have equal sign, it should had been a greater than sign a(t+1)^2+k>a+k I don't know if you caught that or not

OpenStudy (anonymous):

yes, i assumed it was since the problem is dealing with inequality

myininaya (myininaya):

okay cool im still thinking about 3

myininaya (myininaya):

i wonder what happens if we assume f(t)>f(1) and solve for t again im going to try

myininaya (myininaya):

we get -3<t<1

myininaya (myininaya):

you said the last one is true?

myininaya (myininaya):

the interval (-3,1) is not included in the first interval we got so i would think the 3rd one would be false unless i made a mistake somewhere

OpenStudy (anonymous):

right, but possibly the range works somehow? i have the harder time getting why option 3 works...

myininaya (myininaya):

oh wait we have t>1 or t<-3 not -3<t<1

myininaya (myininaya):

oomg wait are you for sure number 1 is right? is it true or false?

myininaya (myininaya):

ok sqrt(x^2)=|x|, so we have sqrt((t+1)^2)>1 and sqrt((t+1)^2))<-1 so we have t>0 or t<-2 and for the last one we got t<-3 or t>1 which are included in that interval so number 3 is true and number 1 is false now what about number 2...

myininaya (myininaya):

for the second one we get -2<t<0 but we said t>0 or t<-2 so number 2 is false

myininaya (myininaya):

i get it. the only true one is number 3 the others are FALSE!

myininaya (myininaya):

what do you think? any questions/

myininaya (myininaya):

i'm fixing to go to bed if you dont have any questions

myininaya (myininaya):

I really don't want to leave if you have questions, but I'm sleepy and I don't know if I should wait or not

OpenStudy (anonymous):

i think i kind of get it, although i'm not sure where the absolute value comes in

myininaya (myininaya):

sqrt(x^2)=|x| just think about it whats sqrt((-3)^2) and is equal to |-3|

myininaya (myininaya):

since I had something square and i took the square root of it that made me think of the absolute value thing where you change the sign of the number when the inequality sign is flipped

OpenStudy (anonymous):

ohh, ok that's right, i think i understand now, that was the one part that i didn't follow

OpenStudy (anonymous):

thank you so much for your help!

myininaya (myininaya):

the first time i did part one i got the interval wrong i don't know if you notice i corrected myself on that one or not

OpenStudy (anonymous):

yes, i saw, thank you for checking

myininaya (myininaya):

we know that interval is true since the question said if it was true right?

myininaya (myininaya):

so we based everything off that result

OpenStudy (anonymous):

exactly, i think we have to assume that base on the problem

myininaya (myininaya):

yes we do because the question says too

myininaya (myininaya):

i really like that problem that was totally awesome goodnight!

OpenStudy (anonymous):

thanks so much for your help, good night to you too!